# Rolling with the Punches: Resilient Contrastive Pre-training under Non-Stationary Drift

### v2

@@ A.3 Contrastive Pre-training
The seminal work of SimCLR (chenSimpleFrameworkContrastive2020) establishes foundational principles for contrastive pretraining through instance discrimination objectives and systematic augmentation strategies. While achieving remarkable performance in static environments, its reliance on fixed augmentation policies and stationary negative sampling assumes temporal consistency of latent patterns – an assumption violated under concept drift scenarios. Nevertheless, SimCLR’s demonstration of invariant representation learning through normalized temperature-scaled loss provides crucial architectural groundwork for developing drift-aware contrastive frameworks, particularly in modeling evolving feature relationships through dynamic positive/negative pair formulation.
@@
Besides, the MoCo framework (heMomentumContrastUnsupervised2020) and its subsequent evolution MoCo v3 (chenEmpiricalStudyTraining2021) establish critical momentum-based mechanisms for contrastive learning through dynamic memory banks and stabilized key encoder updates. It is primarily designed to maintain consistency in negative sample maintenance and gradient stability. However, their fixed momentum schedules and stationary target assumptions limit adaptability to abrupt distribution shifts as we discussed aforementioned. Likewise, BYOL (grillBootstrapYourOwn2020) employs two neural networks—an online network and a target network—that interact and learn from each other. Specifically, the online network is trained to predict the representation of an augmented image view produced by the target network, while the target network is updated using a slow-moving average of the online network’s parameters. Similarly, the SiamSim framework (chenExploringSimpleSiamese2021) advances contrastive representation learning through dynamic similarity calibration, employing a dual-path Siamese architecture with momentum-aligned feature projectors.
@@
Meanwhile, DINO (caronEmergingPropertiesSelfSupervised2021) and DINO v2 (zhouDINOWMWorldModels2024) advance self-distillation paradigms through momentum-based teacher-student mechanisms, leveraging global-local attention consistency in Vision Transformers to learn semantically structured representations. While achieving state-of-the-art in static self-supervised learning, their fixed teacher-student update schedules and monolithic prototype banks implicitly assume stationarity of feature distributions. DINO v2’s introduction of partitioned expertise through specialized sub-networks demonstrates partial adaptability to data variations, suggesting pathways for concept drift mitigation through dynamic expert routing. These works collectively establish critical baselines for stability-aware distillation architectures in non-stationary pre-training scenarios.
@@
Furthermore, DenseCL (wangDenseContrastiveLearning2021a) advances contrastive learning through localized feature alignment, introducing dense instance discrimination that operates at both image-level and pixel-level granularity. By enforcing spatial consistency through region-to-region contrastive pretext tasks, the method enhances model sensitivity to fine-grained visual patterns. While primarily designed for dense prediction tasks, its hierarchical contrast mechanism demonstrates the importance of multi-scale feature stabilization.
@@
Appendix B Experiments
@@ B.1 Datasets
ImageNet-LT (liuLargeScaleLongTailedRecognition2019) is a long-tailed subset derived from the ImageNet (russakovskyImagenetLargeScale2015) dataset, designed to benchmark machine learning models under realistic class imbalance. It contains 115k training images across 1,000 categories, with class frequencies ranging from 1,280 (head classes) to merely 5 (tailed classes), simulating real-world data skewness. The validation set is balanced (20 images per class), while the test set aligns with the standard ImageNet validation data (50,000 images). It serves as a pivotal resource for studying long-tail recognition challenges, including few-shot learning, open-set generalization, and algorithmic fairness, while retaining compatibility with the original ImageNet hierarchy and evaluation protocols.
@@
iNaturalist2018 (vanhornINaturalistSpeciesClassification2018a) is a large-scale dataset designed to study fine-grained recognition under naturally occurring long-tailed class distributions. It comprises 675k training and validation images spanning 5,089 species hierarchically organized into 13 super-categories, with severe imbalance reflecting ecological rarity. The dataset integrates taxonomic metadata, geospatial coordinates, and temporal annotations, supporting research in hierarchical learning, transfer learning, and self-supervised pretraining.
@@
In terms of OOD detection, the Texture (DTD) (cimpoi2014describing) is a comprehensive collection of 5,640 texture images, meticulously organized into 47 categories based on human-centric perceptual attributes. Designed to support texture analysis and recognition tasks, DTD is divided into training, validation, and test sets, each containing 40 images per category. Besides, the iNat-OOD dataset (huangMOSScalingOutofDistribution2021) is a specialized subset of the iNaturalist (vanhornINaturalistSpeciesClassification2018a) dataset, designed for evaluating out-of-distribution (OOD) detection methods in large-scale image classification tasks. The dataset is particularly valuable for testing the robustness of models in identifying novel or unseen categories, especially in ecological and biodiversity contexts. Meanwhile, ImageNet-O (hendrycksNaturalAdversarialExamples2021) is a specialized dataset designed to evaluate the robustness of visual models in detecting out-of-distribution (OOD) samples. It consists of 2,000 images from classes not included in the standard ImageNet-1k (russakovskyImagenetLargeScale2015) dataset, making it a valuable benchmark for testing OOD detection methods. Additionally, OpenImage-O (wangViMOutofDistributionVirtualLogit2022)is a large-scale, manually annotated dataset designed for out-of-distribution (OOD) detection tasks. It is derived from the OpenImage-V3 test set, which contains a diverse and natural distribution of images collected from Flickr without predefined class names or tags. This dataset aims to overcome the limitations of existing OOD benchmarks by providing a more realistic and challenging testbed for evaluating the robustness of computer vision models
@@
Regarding domain shift, ImageNet-V2 (rechtImageNetClassifiersGeneralize2019) is a comprehensive test set designed to evaluate the robustness and generalization of image classification models. It comprises 10,000 images, with 10 images per class, closely following the original labeling protocol of ImageNet. This dataset is instrumental in assessing how well models trained on the original ImageNet dataset can generalize to new, unseen data, making it a valuable resource for advancing computer vision research. The ImageNet-Sketch (wangLearningRobustGlobal2019) is a unique collection of 50,000 hand-drawn sketch images, with 50 images for each of the 1,000 ImageNet classes. Constructed using Google Image queries, this dataset is designed to evaluate models’ ability to learn out-of-domain semantics at the ImageNet scale. ImageNet-R (hendrycksManyFacesRobustness2021) is a collection of 30k images representing 200 ImageNet (russakovskyImagenetLargeScale2015) classes. These images are artistic renditions, including art, cartoons, graffiti, embroidery, origami, paintings, and other forms of creative expressions. The dataset is designed to test the robustness and generalization capabilities of image classification models when faced with non-standard inputs .
B.2 Implementation Details
@@
Table 8: The linear probing hyperparameters.
Downstream Tasks Fine-tuning Linear Probing
Models ViT-S/16 ViT-B/16 ViT-L/16 ViT-S/16 ViT-B/16 ViT-L/16
Training Epochs 100 100 50 90 90 50
Warmup Epochs 5 5 5 10 10 10
Optimizer AdamW AdamW AdamW LARS LARS LARS
Base Learning Rate 5e-4 5e-4 1e-3 0.1 0.1 0.1
Learning Rate Decay Cosine Cosine Cosine Cosine Cosine Cosine
Weight Decay 0.05 0.05 0.05 0.0 0.0 0.0
Eff. Batch Size 1,024 1,024 1,024 8,192 8,192 1,024

### v3

@@ 5.3 Contrastive Pre-training
The seminal work of SimCLR [chenSimpleFrameworkContrastive2020] establishes foundational principles for contrastive pretraining through instance discrimination objectives and systematic augmentation strategies. While achieving remarkable performance in static environments, its reliance on fixed augmentation policies and stationary negative sampling assumes temporal consistency of latent patterns – an assumption violated under concept drift scenarios. Nevertheless, SimCLR’s demonstration of invariant representation learning through normalized temperature-scaled loss provides crucial architectural groundwork for developing drift-aware contrastive frameworks, particularly in modeling evolving feature relationships through dynamic positive/negative pair formulation.
@@
Besides, the MoCo framework [heMomentumContrastUnsupervised2020] and its subsequent evolution MoCo v3 [chenEmpiricalStudyTraining2021] establish critical momentum-based mechanisms for contrastive learning through dynamic memory banks and stabilized key encoder updates. It is primarily designed to maintain consistency in negative sample maintenance and gradient stability. However, their fixed momentum schedules and stationary target assumptions limit adaptability to abrupt distribution shifts as we discussed aforementioned. Likewise, BYOL [grillBootstrapYourOwn2020] employs two neural networks—an online network and a target network—that interact and learn from each other. Specifically, the online network is trained to predict the representation of an augmented image view produced by the target network, while the target network is updated using a slow-moving average of the online network’s parameters. Similarly, the SiamSim framework [chenExploringSimpleSiamese2021] advances contrastive representation learning through dynamic similarity calibration, employing a dual-path Siamese architecture with momentum-aligned feature projectors.
@@
Meanwhile, DINO [caronEmergingPropertiesSelfSupervised2021] and DINO v2 [zhouDINOWMWorldModels2024] advance self-distillation paradigms through momentum-based teacher-student mechanisms, leveraging global-local attention consistency in Vision Transformers to learn semantically structured representations. While achieving state-of-the-art in static self-supervised learning, their fixed teacher-student update schedules and monolithic prototype banks implicitly assume stationarity of feature distributions. DINO v2’s introduction of partitioned expertise through specialized sub-networks demonstrates partial adaptability to data variations, suggesting pathways for concept drift mitigation through dynamic expert routing. These works collectively establish critical baselines for stability-aware distillation architectures in non-stationary pre-training scenarios.
@@
Furthermore, DenseCL [wangDenseContrastiveLearning2021a] advances contrastive learning through localized feature alignment, introducing dense instance discrimination that operates at both image-level and pixel-level granularity. By enforcing spatial consistency through region-to-region contrastive pretext tasks, the method enhances model sensitivity to fine-grained visual patterns. While primarily designed for dense prediction tasks, its hierarchical contrast mechanism demonstrates the importance of multi-scale feature stabilization.
@@
6 Simple and Efficient Implementation with Scalability Advantages
Our RCP method is designed for simplicity and efficiency, integrating seamlessly with existing contrastive learning pipelines. RCP requires only minimal modifications to standard architectures like MoCo v3 [chenEmpiricalStudyTraining2021]. The primary adjustments involve incorporating our intervention module, which operates on the feature embeddings obtained from the student and momentum encoders. To support the larger effective batch sizes beneficial for capturing broader drift adaptation windows, we also leverage a masked contrastive strategy, inspired by recent work [yangMaskedImageContrastive2024, yang2024enhancing].
Furthermore, RCP offers exceptional scalability advantages with respect to both data and model size. Firstly, it robustly accommodates diverse and expanding datasets during pre-training. By mitigating the adverse effects of concept drift, RCP ensures that model performance consistently benefits from increased data volume and variety, aligning with established scaling laws [muennighoffScalingDataConstrainedLanguage2023] without succumbing to drift-induced degradation. Secondly, RCP enhances the training stability of larger models in dynamic data environments. This improved stability facilitates the pre-training of models with significantly more parameters and deeper architectures, pushing the boundaries of what can be learned from evolving data streams.
7 Experiments
@@ 7.1 Datasets
ImageNet-LT [liuLargeScaleLongTailedRecognition2019] is a long-tailed subset derived from the ImageNet [russakovskyImagenetLargeScale2015] dataset, designed to benchmark machine learning models under realistic class imbalance. It contains 115k training images across 1,000 categories, with class frequencies ranging from 1,280 (head classes) to merely 5 (tailed classes), simulating real-world data skewness. The validation set is balanced (20 images per class), while the test set aligns with the standard ImageNet validation data (50,000 images). It serves as a pivotal resource for studying long-tail recognition challenges, including few-shot learning, open-set generalization, and algorithmic fairness, while retaining compatibility with the original ImageNet hierarchy and evaluation protocols.
@@
iNaturalist2018 [vanhornINaturalistSpeciesClassification2018a] is a large-scale dataset designed to study fine-grained recognition under naturally occurring long-tailed class distributions. It comprises 675k training and validation images spanning 5,089 species hierarchically organized into 13 super-categories, with severe imbalance reflecting ecological rarity. The dataset integrates taxonomic metadata, geospatial coordinates, and temporal annotations, supporting research in hierarchical learning, transfer learning, and self-supervised pretraining.
@@
In terms of OOD detection, the Texture (DTD) [cimpoi2014describing] is a comprehensive collection of 5,640 texture images, meticulously organized into 47 categories based on human-centric perceptual attributes. Designed to support texture analysis and recognition tasks, DTD is divided into training, validation, and test sets, each containing 40 images per category. Besides, the iNat-OOD dataset [huangMOSScalingOutofDistribution2021] is a specialized subset of the iNaturalist [vanhornINaturalistSpeciesClassification2018a] dataset, designed for evaluating out-of-distribution (OOD) detection methods in large-scale image classification tasks. The dataset is particularly valuable for testing the robustness of models in identifying novel or unseen categories, especially in ecological and biodiversity contexts. Meanwhile, ImageNet-O [hendrycksNaturalAdversarialExamples2021] is a specialized dataset designed to evaluate the robustness of visual models in detecting out-of-distribution (OOD) samples. It consists of 2,000 images from classes not included in the standard ImageNet-1k [russakovskyImagenetLargeScale2015] dataset, making it a valuable benchmark for testing OOD detection methods. Additionally, OpenImage-O [wangViMOutofDistributionVirtualLogit2022]is a large-scale, manually annotated dataset designed for out-of-distribution (OOD) detection tasks. It is derived from the OpenImage-V3 test set, which contains a diverse and natural distribution of images collected from Flickr without predefined class names or tags. This dataset aims to overcome the limitations of existing OOD benchmarks by providing a more realistic and challenging testbed for evaluating the robustness of computer vision models
@@
Regarding domain shift, ImageNet-V2 [rechtImageNetClassifiersGeneralize2019] is a comprehensive test set designed to evaluate the robustness and generalization of image classification models. It comprises 10,000 images, with 10 images per class, closely following the original labeling protocol of ImageNet. This dataset is instrumental in assessing how well models trained on the original ImageNet dataset can generalize to new, unseen data, making it a valuable resource for advancing computer vision research. The ImageNet-Sketch [wangLearningRobustGlobal2019] is a unique collection of 50,000 hand-drawn sketch images, with 50 images for each of the 1,000 ImageNet classes. Constructed using Google Image queries, this dataset is designed to evaluate models’ ability to learn out-of-domain semantics at the ImageNet scale. ImageNet-R [hendrycksManyFacesRobustness2021] is a collection of 30k images representing 200 ImageNet [russakovskyImagenetLargeScale2015] classes. These images are artistic renditions, including art, cartoons, graffiti, embroidery, origami, paintings, and other forms of creative expressions. The dataset is designed to test the robustness and generalization capabilities of image classification models when faced with non-standard inputs.
@@
Table 9: The hyperparameters of fine-tuning and linear probing.
Downstream Tasks Fine-tuning
Models ViT-S/16 ViT-B/16 ViT-L/16
Training Epochs 100 100 50
Warmup Epochs 5 5 5
Optimizer AdamW AdamW AdamW
Base Learning Rate 5e-4 5e-4 1e-3
Learning Rate Decay Cosine Cosine Cosine
Weight Decay 0.05 0.05 0.05
Eff. Batch Size 1,024 1,024 1,024
Downstream Tasks Linear Probing
Models ViT-S/16 ViT-B/16 ViT-L/16
Training Epochs 90 90 50
Warmup Epochs 10 10 10
Optimizer LARS LARS LARS
Base Learning Rate 0.1 0.1 0.1
Learning Rate Decay Cosine Cosine Cosine
Weight Decay 0.0 0.0 0.0
Eff. Batch Size 8,192 8,192 1,024
7.2 Implementation Details
@@
7.3 Additional Experiments
Additionally, we also conducted zero-shot experiments on TIC-DataComp(L) to verify our methods, following the settings of TiC-CLIP[garg2024ticclip]. Our methods achieve superior results compared to other methods on TiC-CLIP datasets. Furthermore, it is noted that compared methods are based on vision-language models with image-text pairs, while we only used images for testing ViT.
ImageNet
ImageNet
dist shift
Flickr30k
Sequential 44.7 37.4 48.4
Patching 45.8 38.9 49.7
Cumulative-Exp 47.3 39.6 50.8
Cumulative-Equal 47.7 40.3 51.8
Cumulative-All 53.0 44.3 54.4
Oracle 53.6 44 53.9
[…]
Table 10: Zero-shot experiments on TIC-DataComp(L).
Moreover, we have added more experimental results of DINO with our proposed module to demonstrate the generalization of our module on various contrastive frameworks. The results show that our proposed module can also improve the performance of DINO under non-stationary drift. It demonstrates that our proposed intervention itself is agnostic to the backbone or the specific contrastive framework: any model that maintains a momentum-updated teacher and processes mini-batches whose distribution may drift can incorporate our proposed module.
Many Medium Few All
ImageNet-LT
DINO 64.8 35.5 11.3 43.3
DINO+Ours 67.1 37.7 14.9 45.9
MoCo v3 70.8 40.7 14.3 48.5
MoCo v3+Ours 72.2 43.1 16 50.3
iNaturalist2018
DINO 76.7 67.1 61.2 65.5
DINO+Ours 77.1 67.5 62.1 66.1
MoCo v3 76.6 65.8 62.8 65.6
MoCo v3+Ours 77.7 68.7 63.8 67.5
Table 11: Generalization of our module on various contrastive frameworks.